OpenStudy (anonymous):

How would you solve: 1/2 + 1/3 = 1/x ? Can anyone help guide me through it so I can figure out how to do the others? Thanks.

OpenStudy (dls):

\[\Huge \frac{1}{2}+\frac{1}{3}=\frac{1}{x}\] Solve LHS by taking LCM.

mathslover (mathslover):

Take LCM of the denominators .

OpenStudy (anonymous):

6x right?

mathslover (mathslover):

No.

mathslover (mathslover):

LCM of 2 and 3 = 6 only

mathslover (mathslover):

Now, multiply the LCM of 2 and 3 both sides : \((\cfrac{1}{2} +\cfrac{1}{3})\times \color{blue}{6} = \cfrac{1}{x} \times \color{blue}{ 6}\)

OpenStudy (anonymous):

oh ok I see so once we multiply LCM we would have 3/6 + 2/6 = 1/x which equals 5/6 = 1/x right?

OpenStudy (dls):

1/x=5/6 x=?

OpenStudy (anonymous):

Thanks once u cross multiply and solve for x it's x=6/5 :) Thanks 4 guiding me guys :D

mathslover (mathslover):

Yes x = 6/5